OpenStudy (anonymous):

given that triangle abc is congruent to triangle def, ac = 21x - 19, and df = 3x + 17, find the measures of segment ac a. 2 b. 9 c. 21 d. 23

jimthompson5910 (jim_thompson5910):

place ABC directly over DEF 

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

"given that triangle abc is congruent to triangle def" and using CPCTC, we know that AC = DF

jimthompson5910 (jim_thompson5910):

AC = DF 21x - 19 = 3x + 17 solve for x once you have the value of x, you use it to find the length of AC

OpenStudy (anonymous):

i cant get x ); i kkeep getting random numbers plz help out

OpenStudy (anonymous):

its 2 (: wat i do next????

OpenStudy (anonymous):

is 2 my final answer??

jimthompson5910 (jim_thompson5910):

x = 2 is not your final answer

jimthompson5910 (jim_thompson5910):

plug x = 2 into AC = 21x-19 to compute the length of AC

Directrix (directrix):

From JThompson's work: AC = DF 21x - 19 = 3x + 17 You got x = 2

Directrix (directrix):

AC = 21x-19 AC = 21*2 - 19 AC = ? @mikepeters1994

Directrix (directrix):

@mikepeters1994 Please stay in the thread while we are actively working. Thanks.

OpenStudy (anonymous):

ok thx